Assuming the standard deviation is unchanged and that the usage is normally distributed, provide an expression for calculating a 99% confidence interval for the mean usage in the March quarter of 2006. (Exactly one option must be correct)



1
Expert's answer
2021-05-07T10:28:32-0400

Since "\u03c3" is assumed known,


we use the interval "x̄ ± z \dfrac{σ}{ n},"


 where "\u03c3 = 81, n = 30" and where z is chosen to ensure that "P(|Z|\u2264 z) = 0.99."


From the normal tables, "P(|Z|\u2264 2.575) = 0.99" (because "P(Z < 2.575) = 0.995)" and so we use "z = 2.575."


99% confidence interval expression "= \bar{x}\pm 2.575\times \dfrac{\sigma}{n}"